# Collaborative Rational Speech Act: Pragmatic Reasoning for Multi-Turn Dialog

### v1

@@ 5.2 Numerical results and discussion
Figure 2 presents the performance of the CRSA model compared to baseline models for $\alpha=2.5$. Each curve corresponds to a different model evaluated over 500 rounds of the game. The top plot displays task accuracy, measured as the proportion of correct guesses obtained by taking the argmax of the listener’s posterior probability. As accuracy may not be fully representative of the confidence on the decision made by the listener, we also show the Information Gain (in the bottom plot) for each turn $t$, computed as the difference $\textrm{IG}(L_{t})=H_{P}(Y|M_{L_{t}})-H_{L}(Y|U_{t},M_{L_{t}},W_{t})$. That is, given a set of $N$ rounds (all with the same number of turns), the listener’s conditional entropy is defined as $H_{L}(Y|U_{t},M_{L_{t}},W_{t})=-1/N\sum_{i=1}^{N}\log L(y^{(i)}|u^{(i)}_{t},w^{(i)}_{t},m^{(i)}_{L_{t}})$, and the conditional entropy of the prior is defined as $H_{P}(Y|M_{L_{t}})=-1/N\sum_{i=1}^{N}\log P(y^{(i)}|m^{(i)}_{L_{t}})$, where the super-index $(i)$ denotes the value at round $i$. As $P(y|m_{L_{t}})$ takes no account for the interchanged utterances, this metric could be interpreted as the amount of information gained by using the utterances of the dialog up to turn $t$. For all models where there is iteration, we run the model until the gain converged using a tolerance of $31e-$ 3, so the number of iterations may vary between each turn. We tried various values of $\alpha>1$ and all values showed best performance of the CRSA model. For values $\alpha\leq 1$, all iterative algorithms always produced uniform distributions.
@@
As shown in the plots, the CRSA model outperforms all baselines across both metrics. Moreover, incorporating a lexicon that depends on the past $w_{t}$ neither improves nor diminishes performance, suggesting that the information encoded in Eq. (10) is already effectively captured by the CRSA model. In contrast, the information in Eq. 10 is not captured by the YRSA-$W_{t}$ model, which appears to improve as the conversation progresses. As expected, models that do not incorporate dialog history maintain consistent performance across turns, with variations driven only by role changes. We also observed that the CRSA model’s variance decreases over time, although this is not shown in the plots for clarity.
@@
Figure 3 presents an example of a dialogue between the agents, along with their internal belief states at each turn. This dialog was generated by sampling the pragmatic speaker distribution at each turn. Each column displays the value of $B_{S,t}(m_{L_{t}})$ for each possible meaning $m_{L_{t}}$ of the listener at turn $t$. Notably, as the conversation progresses, the meanings associated with previously uttered messages tend to gain higher belief values, reflecting a refinement in the speaker’s inference about the listener’s state. We note that speaker at turn 5 produce the "Position 2" utterance, which is a little uninintuitive. However, since utterances are sampled from the full pragmatic speaker distribution rather than chosen greedily and "Position 2" has non-zero probability, this utterance was drawn randomly out of other possible utterances. Such decoding strategy can be interpreted as an exploratory strategy. Importantly, Agent B’s return to "Position 5" in the following turn is consistent with its high posterior belief. We also note that the value that maximizes $B_{S,t}(m_{L_{t}})$ at turn 6 does not correspond exactly to the correct meaning, but it is a close approximation since the utterance “Position 6” never occurred during the round. This supports interpreting $B_{S,t}(m_{L_{t}})$ as the speaker’s belief about the listener’s meaning $m_{L_{t}}$ at turn $t$.
@@ 6 Modeling Conversations Using Pragmatic LLMs
In this Section, we provide preliminary evidence that the CRSA model can produce reasonably good estimations of both the likelihood of each utterance $u_{t}$ and the target $y$ of the task in doctor–patient conversations, which is the disease corresponding to the symptoms described by the patient. To this end, we used the MDDial dataset (Macherla et al., 2023), which consists of template-based conversations between a doctor and a patient. In each dialog, the patient is assigned a subset of predefined symptoms, and the doctor must determine the correct disease from a set of possible pathologies.
@@
As anticipated in Section 4.2.2, in order to apply the pragmatic models, we compute the literal speaker with equation 8 using a pre-trained LLaMA3.2–1B-Instruct language model. In this equation, $\mathrm{prompt}(m_{S_{t}})$ is the text used to prompt the model with the relevant medical scenario. When $S_{t}$ is the doctor, the prompt includes specific instructions to ask questions and produce a diagnosis, followed by two example doctor–patient conversations. When $S_{t}$ is the patient, the prompt instructs the model to play the role of the patient. It uses the same conversation examples as in the doctor prompt but additionally includes the patient’s current symptoms at that turn. The full prompts used can be found in Appendix C. In order to save computation, we pre-computed the value of the literal speaker for each possible combination of utterance $u_{t}$ and symptoms $m_{S_{t}}$ for each turn $t$ and then applied the update equations to that.
Finally, since each dialog ends with the doctor explicitly stating the disease, we compute the final listener distribution $L_{T^{(i)}}(y|u_{T^{(i)}},m_{L_{T^{(i)}}},w_{T^{(i)}})$ for the last turn ${T^{(i)}}$ of round $i$ by replacing the disease name in the utterance with each candidate diagnosis $y\in\mathcal{Y}$. This probability is used at each round as the listener probability of the literal model.
6.1 Numerical results and discussion
@@
To evaluate performance, we compute the speaker entropy as $H_{S}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T_{i}}\log S_{t}\big{(}u^{(i)}_{t}\mid w^{(i)}_{t},m^{(i)}_{S_{t}}\big{)},$ where $N$ is the number of rounds and $T_{i}$ is the number of turns in round $i$. For the Listener, we compute $H_{L}=-\frac{1}{N}\sum_{i=1}^{N}L_{T_{i}}\big{(}y\mid u_{T_{i}},m_{L_{T_{i}}},w_{T_{i}}\big{)},$ which is analogous to the method described in Section 5, but considers only the distribution at the final turn of each round.
The results are presented in Table 1 for 65 samples of the dataset and for a value of $\alpha=2.5$, which is the same as used for Section 5. We observed the same trend mentioned in that Section when varying the value of $\alpha$. The CRSA model achieves substantially lower entropy for both the speaker and the listener, outperforming both the RSA and Literal baselines. While the utterances in the MDDial dataset are not naturally produced—since they follow pre-defined templates—these results offer initial evidence that CRSA captures key aspects of pragmatic reasoning and task-oriented belief updates across turns.
$H_{S}$ $H_{L}$
CRSA 8.18 2.19
RSA 14.27 2.86
Literal 18.00 3.29
@@
Table 1: Entropies for $\alpha=2.5$ of the listener and the speaker for each model, computed for 65 samples of the MDDial dataset.
@@ 7 Possible Future Directions of this work
Finally, there are many ways in which the original gain function from which the equations of the model are derived could be modified depending on the application scenario. For instance, situations in which the meanings are not fixed in time, or where more than two agents participate in the dialogue, can also fit within a similar procedure to that used in this work. This allows for the introduction of pragmatic reasoning in more realistic scenarios in the same mathematically grounded way as was done in the current work.
@@ 8 Summary and Concluding Remarks
In this work, we introduced the Collaborative Rational Speech Act (CRSA) framework, an information-theoretic extension of RSA tailored for principled pragmatic reasoning in multi-turn, task-oriented dialogues. By integrating a novel multi-turn gain function grounded in interactive rate-distortion theory, CRSA effectively models the evolving belief dynamics of both interlocutors, overcoming key limitations of traditional RSA in collaborative contexts. Our preliminary results demonstrate that CRSA successfully captures the progression of shared understanding, partner beliefs, and utterance generation, providing the way for more natural and efficient communication in complex conversational settings.
CRSA lays the foundation for developing conversational agents driven by mathematically grounded principles of pragmatic reasoning. This principled formulation enhances both the interpretability and controllability of agent behavior, enabling the construction of language models that move beyond surface-level fluency to demonstrate structured, socially coherent, and contextually appropriate dialogue. In this way, CRSA represents a significant step toward building pragmatic agents whose interactions are not only effective but also firmly rooted in the formal theory of communication.
@@
This work focuses on simulated referential games and template-based doctor–patient dialogues, which, while controlled and insightful, do not capture the full variability and complexity of real-world conversations. Additionally, the CRSA framework relies on a fixed, predefined set of possible utterances at each turn, limiting its applicability to open-ended or generative dialogue scenarios involving variable-length token sequences. These factors currently restrict the scalability of our approach to more naturalistic domains. Future work will aim to overcome these limitations by extending CRSA to handle dynamically generated utterance spaces and by evaluating its effectiveness in less structured, real-world conversational settings.
@@
This work presents a theoretically grounded framework for pragmatic reasoning in multi-turn dialogs. It is primarily methodological and does not involve direct deployment or interaction with real users. The datasets employed—simulated referential games and template-based medical dialogues—are synthetic and contain no personal or sensitive data.
@@ Appendix A Detailed Expressions of the YRSA Model
If the same procedure is applied to the gain of equation 4.1 (the one corresponding to the YRSA model), then the following equations are obtained:
@@ Appendix C Prompts used in the MDDial dataset
You are an assistant that simulates to be a patient who has a disease and describes the symptoms to the user, which is a medical doctor.
Here is an example of a conversation between the assitant (i.e., the patient) and the user (i.e., the doctor). You are experiencing the following symptoms: Fear of cold, Poor sleep, Feel sick and vomit, Syncope, Increased stool frequency, Blood in the tears, Hoarse, Loss of appetite, Dizzy, Expectoration, Pharynx discomfort, Limb numbness, Acid reflux, Thirst, Diarrhea, Difficulty breathing, Stuffy nose, Vomiting, Bitter, Runny nose, Bloody stools, Bloating, Fatigue, Hard to swallow, Chest tightness, Shortness of breath, Palpitations, Fever, Headache, Constipation, Thin, Dizziness, Body aches, Diplopia, Hemoptysis, Burning sensation behind the breastbone, Sweating, Hiccup, Poor spirits, Frequent urination, Black stool, Consciousness disorder, Thin white moss, Edema, Anorexia, Cough, Stomach ache, Pain behind the breastbone, Sleep disorder, Hematemesis, Nausea, Twitch, Hiccough, Chest tightness and shortness of breath
Assistant: I have been feeling Burning sensation behind the breastbone
User: Oh, do you have any Stomach ache?
Assistant: I am experiencing that sometimes
User: Oh, do you have any Acid reflux?
Assistant: Yes Doctor, I am feeling that as well
@@
Here is an example of a conversation between the assitant (i.e., the patient) and the user (i.e., the doctor). You are experiencing the following symptoms: Pain in front of neck, Nose bleeding, Fear of cold, Syncope, Feel sick and vomit, Poor sleep, Headache and dizziness, Hazy, Loss of appetite, Hearing loss, Dizzy, Cry, Diarrhea, Difficulty breathing, Stuffy nose, Vomiting, Runny nose, Bloating, Fatigue, Palpitations, Chest tightness, Headache, Fever, Dizziness, Incontinence, Poor spirits, Redness, Waist pain, Unconsciousness, Vertigo, Consciousness disorder, Head trauma pain, Poor physical activity, Anorexia, Cough, Stomach ache, Pain behind the breastbone, Hematemesis, Sleep disorder, Earache, Nausea, Tinnitus, Twitch, Limb numbness, Chest tightness and shortness of breath
Assistant: I have Nausea and Dizziness
User: I believe you are having from Traumatic brain injury.
Now, participate in a real conversation with the user. You are experiencing the following symptoms: {patient symptoms}
@@
You are an assistant that simulates to be a doctor who is diagnosing a patient based on the symptoms that he or she describes. You can ask questions to the patient, but ultimately, you have to provide a diagnosis based on the symptoms described by the patient.
Here is an example of a conversation between the assitant (i.e., the doctor) and the user (i.e., the patient):
User: I have been feeling Burning sensation behind the breastbone
Assistant: Oh, do you have any Stomach ache?
User: I am experiencing that sometimes
Assistant: Oh, do you have any Acid reflux?
User: Yes Doctor, I am feeling that as well
Assistant: This could probably be Esophagitis.
Here is an example of a conversation between the assitant (i.e., the doctor) and the user (i.e., the patient):
User: I have Nausea and Dizziness
Assistant: I believe you are having from Traumatic brain injury.
Now, participate in a real conversation with the user. You can ask questions to the patient, but ultimately, you have provide a diagnosis based on the symptoms described by the patient.

### v2

@@ 5.2 Numerical results and discussion
Figure 2 presents the performance of the CRSA model compared to baseline models for $\alpha=2.5$. Each curve corresponds to a different model evaluated over 500 rounds of the game. The top plot displays task accuracy, measured as the proportion of correct guesses obtained by taking the argmax of the listener’s posterior probability. As accuracy may not be fully representative of the confidence on the decision made by the listener, we also show the Information Gain (in the bottom plot) for each turn $t$, computed as the difference $\textrm{IG}(L_{t})=H_{P}(Y|M_{L_{t}})-H_{L}(Y|U_{t},M_{L_{t}},W_{t})$. That is, given a set of $N$ rounds (all with the same number of turns), the listener’s conditional entropy is defined as $H_{L}(Y|U_{t},M_{L_{t}},W_{t})=-1/N\sum_{i=1}^{N}\log L_{t}(y^{(i)}|u^{(i)}_{t},w^{(i)}_{t},m^{(i)}_{L_{t}})$, and the conditional entropy of the prior is defined as $H_{P}(Y|M_{L_{t}})=-1/N\sum_{i=1}^{N}\log P(y^{(i)}|m^{(i)}_{L_{t}})$, where the super-index $(i)$ denotes the value at round $i$. As $P(y|m_{L_{t}})$ takes no account for the interchanged utterances, this metric could be interpreted as the amount of information gained by using the utterances of the dialog up to turn $t$. For all models where there is iteration, we run the model until the gain converged using a tolerance of 1e31e-3, so the number of iterations may vary between each turn. We tried various values of $\alpha>1$ and all values showed best performance of the CRSA model. For values $\alpha\leq 1$, all iterative algorithms always produced uniform distributions.
@@
As shown in the plots, the CRSA model outperforms all baselines across both metrics. Moreover, incorporating a lexicon that depends on the past $w_{t}$ neither improves nor diminishes performance, suggesting that the information encoded in Equation (10) is already effectively captured by the CRSA model. In contrast, the information in Equation (10) is not captured by the YRSA-$W_{t}$ model, which appears to improve as the conversation progresses. As expected, models that do not incorporate dialog history maintain consistent performance across turns, with variations driven only by role changes. We also observed that the CRSA model’s variance decreases over time, although this is not shown in the plots for clarity.
@@
Figure 3 presents an example of a dialog between the agents, along with their internal belief states at each turn. This dialog was generated by sampling the pragmatic speaker distribution at each turn. Each column displays the value of $B_{S,t}(m_{L_{t}})$ for each possible meaning $m_{L_{t}}$ of the listener at turn $t$. Notably, as the conversation progresses, the meanings associated with previously uttered messages tend to gain higher belief values, reflecting a refinement in the speaker’s inference about the listener’s state. We note that speaker at turn 5 produce the "Position 2" utterance, which is a little uninintuitive. However, since utterances are sampled from the full pragmatic speaker distribution rather than chosen greedily and "Position 2" has non-zero probability, this utterance was drawn randomly out of other possible utterances. Such decoding strategy can be interpreted as an exploratory strategy. Importantly, Agent B’s return to "Position 5" in the following turn is consistent with its high posterior belief. We also note that the value that maximizes $B_{S,t}(m_{L_{t}})$ at turn 6 does not correspond exactly to the correct meaning, but it is a close approximation since the utterance “Position 6” never occurred during the round. This supports interpreting $B_{S,t}(m_{L_{t}})$ as the speaker’s belief about the listener’s meaning $m_{L_{t}}$ at turn $t$.
@@ 6 Modeling Conversations Using Pragmatic LLMs
In this section, we present preliminary evidence that the CRSA model can estimate both utterance likelihoods and task targets in doctor–patient conversations. Specifically, it improves the mean perplexity for conversation utterances and of the final diagnosis prediction, compared to the raw outputs of the LLM. To this end, we used the MDDial dataset (Macherla et al., 2023), which consists of template-based conversations between a doctor and a patient. In each dialog, the patient is assigned a subset of predefined symptoms, and the doctor must determine the correct disease from a set of possible pathologies.
Methodology
@@
As anticipated in Section 4.2.1, in order to apply the pragmatic models, we compute the literal speaker with equation (8) using a pre-trained LLaMA3.2–1B-Instruct [^5] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct language model. In this equation, $\mathrm{prompt}(m_{S_{t}})$ is the text used to prompt the model with the relevant medical scenario. When $S_{t}$ is the doctor, the prompt includes specific instructions to ask questions and produce a diagnosis, followed by two example doctor–patient conversations. When $S_{t}$ is the patient, the prompt instructs the model to play the role of the patient. It uses the same conversation examples as in the doctor prompt but additionally includes the patient’s current symptoms at that turn. The full prompts used can be found in Appendix C. Importantly, we assume that the set of possible utterances is pre-defined, and we compute the speaker probability over this set. We use the literal speaker as lexicon in Equation (4.2.1) for computing the literal listener.
We compute $P(m_{patient},m_{doctor},y_{diagnose})$ by counting the number of times that symptoms uttered by the patient appear in the context of a certain diagnosis. Note that for this case, the number of possible meanings for the doctor is 1, since it is assumed he/she has allways the same background on knowledge in the field.
Metrics
@@
To evaluate performance, we compute the speaker perplexity as $PPL=\frac{1}{N}\sum_{i=1}^{N}\exp\left(-\sum_{t=1}^{T_{i}}\log S_{t}\big(u^{(i)}_{t}\mid w^{(i)}_{t},m^{(i)}_{S_{t}}\big)\right)$, where $N$ is the number of rounds and $T_{i}$ is the number of turns in round $i$. For the listener, we compute the task success rate using the listener of the last step $T_{i}$ of each round $TSR=-\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}_{\{y^{(i)}\,=\,\operatorname*{arg\,max}_{y\in\mathcal{Y}}L_{T_{i}}\big(y\mid u_{T_{i}},m_{L_{T_{i}}},w_{T_{i}}\big)\}}$ , which is analogous to the method described in Section 5.
Results
The results are presented in Table 1 for the train split of the dataset (1878 samples) and for a value of $\alpha=2.5$, which is the same as used for Section 5. We observed the same trend mentioned in that Section when varying the value of $\alpha$. The CRSA model achieves best performance in terms of both perplexity and accuracy (task sucess rate) compared to the classic RSA and the Literal models. We noted however that the difference between the classic RSA and CRSA is not very large, possibly due to the fact that for this task the agent playing the role of the doctor struggles to obtain a good estimation of the patient’s belief. Since this estimation is close to uniform at all steps, both models converge to nearly identical equations and thus exhibit similar performance. Still, the low task success rate suggests that the setting is inherently difficult and may require more discriminative models. CRSA provides a useful structure in this direction, and it is plausible that coupling it with a stronger LLM could yield a more informative initialization of the speaker and, in turn, improved task performance.
Speaker PPL Task success rate
CRSA 15.362 0.085
RSA 15.395 0.085
Literal 60.835 0.079
@@ Results
Table 1: Speaker perplexity (PPL) and Task success rate for $\alpha=2.5$ of the listener and the speaker for each model, computed for the MDDial dataset.
@@ 7 Possible Future Directions of this work
Finally, there are many ways in which the original gain function from which the equations of the model are derived could be modified depending on the application scenario. For instance, situations in which the meanings are not fixed in time, or where more than two agents participate in the dialog, can also fit within a similar procedure to that used in this work. This allows for the introduction of pragmatic reasoning in more realistic scenarios in the same mathematically grounded way as was done in the current work.
@@ 8 Summary and Concluding Remarks
In this work, we introduced the Collaborative Rational Speech Act (CRSA) framework, an information-theoretic extension of RSA tailored for principled pragmatic reasoning in multi-turn, task-oriented dialogs. By integrating a novel multi-turn gain function grounded in interactive rate-distortion theory, CRSA effectively models the evolving belief dynamics of both interlocutors, overcoming key limitations of traditional RSA in collaborative contexts. Our preliminary results demonstrate that CRSA successfully captures the progression of shared understanding, partner beliefs, and utterance generation, providing the way for more natural and efficient communication in complex conversational settings.
CRSA lays the foundation for developing conversational agents driven by mathematically grounded principles of pragmatic reasoning. This principled formulation enhances both the interpretability and controllability of agent behavior, enabling the construction of language models that move beyond surface-level fluency to demonstrate structured, socially coherent, and contextually appropriate dialog. In this way, CRSA represents a significant step toward building pragmatic agents whose interactions are not only effective but also firmly rooted in the formal theory of communication.
@@
This work focuses on simulated referential games and template-based doctor–patient dialogs, which, while controlled and insightful, do not capture the full variability and complexity of real-world conversations. Additionally, the CRSA framework relies on a fixed, predefined set of possible utterances at each turn, limiting its applicability to open-ended or generative dialog scenarios involving variable-length token sequences. These factors currently restrict the scalability of our approach to more naturalistic domains. Future work will aim to overcome these limitations by extending CRSA to handle dynamically generated utterance spaces and by evaluating its effectiveness in less structured, real-world conversational settings.
@@
This work presents a theoretically grounded framework for pragmatic reasoning in multi-turn dialogs. It is primarily methodological and does not involve direct deployment or interaction with real users. The datasets employed—simulated referential games and template-based medical dialogs—are synthetic and contain no personal or sensitive data.
@@
A final potential ethical concern is the risk of anthropomorphizing AI systems when they are described as communicative agents. While the agent metaphor is useful for modeling and analysis, it may inadvertently suggest that such systems possess autonomy, intentionality, or even consciousness. We stress that this is not the case: our use of agent-like terminology is strictly metaphorical and does not imply any deep philosophical claims about the nature of AI systems.
Acknowledgments
We thank the anonymous reviewers for their insightful feedback. This work benefited from the resources (GPUs and CPUs) of Lab-AI, an institution member of Université Paris- Saclay, for running the experiments.
@@ Appendix A Detailed Expressions of the YRSA Model
If the same procedure is applied to the gain of Equation (4.1) (the one corresponding to the YRSA model), then the following equations are obtained:
@@ Appendix C Prompts used in the MDDial dataset
You are an assistant that simulates to be
a patient who has a disease and describes
the symptoms to the user, which is a
medical doctor.
Here is an example of a conversation
between the assitant (i.e., the patient)
and the user (i.e., the doctor). You are
experiencing the following symptoms:
Acid reflux, Stomach ache, Bloating
Assistant: Hi Doctor, I am having Acid
reflux
User: In that case, do you have any
Stomach ache?
Assistant: Yes most of the times
User: In that case, do you have any
Bloating?
Assistant: Yes most of the times
@@
Here is an example of a conversation
between the assitant (i.e., the patient)
and the user (i.e., the doctor). You are
experiencing the following symptoms:
Eye swelling
Assistant: Recently, I am experiencing
Eye swelling
User: Is it? Then do you experience Cry?
Assistant: No, I never had anything like
that.
User: Oh, do you have any Fever?
Assistant: No, I never had anything like
that.
User: Oh, do you have any Photophobia?
Assistant: No, I don’t have that
User: Ok, this means you might be having
Conjunctivitis.
Now, participate in a real conversation
with the user. You are experiencing the
following symptoms:
{patient symptoms}
@@
You are an assistant that simulates to be
a doctor who is diagnosing a patient based
on the symptoms that he or she describes.
You can ask questions to the patient, but
ultimately, you have to provide a diagnosis
based on the symptoms described by the
patient.
Here is an example of a conversation
between the assitant (i.e., the doctor)
and the user (i.e., the patient). The
patient is experiencing the following
symptoms:
User: Hi Doctor, I am having Acid reflux
Assistant: In that case, do you have
any Stomach ache?
User: Yes most of the times
Assistant: In that case, do you have
any Bloating?
User: Yes most of the times
Assistant: This could probably be
Esophagitis.
Here is an example of a conversation
between the assitant (i.e., the doctor)
and the user (i.e., the patient). The
patient is experiencing the following
symptoms:
User: Recently, I am experiencing Eye
swelling
Assistant: Is it? Then do you
experience Cry?
User: No, I never had anything like
that.
Assistant: Oh, do you have any Fever?
User: No, I never had anything like
that.
Assistant: Oh, do you have any
Photophobia?
User: No, I don’t have that
Assistant: Ok, this means you might
be having Conjunctivitis.
Now, participate in a real conversation
with the user. You can ask questions to
the patient, but ultimately, you have
to provide a diagnosis based on the
symptoms described by the patient.